If the tangents at two points of a parabola are at right angles, then show that they intersects at a point on the directrix.
1
Expert's answer
2020-11-16T19:26:17-0500

Let equation of the parabola is, "y^2=4ax"

Tangent at any point "(x_1,y_1)" is given by "yy_1 = 2a(x+x_1)"


Let any two points on the parabola be "(at_1^2,2at_1),(at_2^2,2at_2)"

Then equation of the tangent at "(at_1^2,2at_1)" is given by,

"2at_1y=2a(x+at_1^2)"

"t_1y=(x+at_1^2) \implies y=\frac{1}{t_1}(x+at_1^2)"


Equation of the tangent at point "(at_2^2,2at_2)" is given by

"2at_2y=2a(x+at_2^2)"

"t_2y=(x+at_2^2) \implies y=\frac{1}{t_2}(x+at_2^2)"


Since, these these tangents are perpendicular then,

"\frac{1}{t_1}\frac{1}{t_2} = -1 \implies \frac{1}{t_2} = -t_1" (1)


Also these tangents intersect, then point of intersection is

"\frac{1}{t_1}(x+at_1^2)=\frac{1}{t_2}(x+at_2^2)"

from equation (1), I can write,


"(x+at_1^2) = -t_1^2(x+a\frac{1}{t_1^2})"


"(1-t_1^2)x = -a(1-t_1^2)"


"\implies x=-a"

It is equation of the directrix.